OpenStudy (anonymous):

An 80.0-kg astronaut is working on the engines of his ship, which is drifting through space with a constant velocity. The astronaut, wishing to get a better view of the Universe, pushes against the ship and much later finds himself 30.0 m behind the ship and at rest with respect to it. Without a thruster, the only way to return to the ship is to throw his 0.500-kg wrench directly away from the ship. If he throws the wrench with a speed of 20.0 m/s relative to the ship, how long does it take the astronaut to reach the ship?

OpenStudy (anonymous):

\[M_{wrench} \times Speed_{wrench}+M_{astronaut} \times Speed_{astronaut}=0\] With this you can calculate the speed of the astronaut versus the ship once he has thrown the wrench. The negative value you get meaning that the velocity of the astronaut has opposite direction to the velocity of the wrench, which means the astronaut starts to move towards the ship. Anyway, I would not recommend this method to NASA

OpenStudy (anonymous):

Basically because in outer space there are no hardware stores where astronauts may get a spare wrench

OpenStudy (anonymous):

wow thats great! Thanks!

OpenStudy (anonymous):

Let x m/s be the final speed of astronaut. Then, by conservation of momentum, 80x - .5*20 = 0 (Assuming the direction of movement of wrench as negative) x = 1/8 m/s Hence, to cover a distance of 30 m, time taken = distance/speed = 30*8 = 240 s.